Circular No. 6591

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) G. Narayanan, Steward Observatory; H. M. Butner, Submillimeter

Telescope Observatory (SMTO); J. McMullin, National Radio Astronomical Observatories; and D. Muders, SMTO, report observations of HCO^{+} and CO with the 10-m Heinrich Hertz Telescope on Mar. 5 and 9: “The J(3–2) line of HCO^{+} at 267.6 GHz was detected on Mar. 5 with a mainbeam temperature of 0.94 ± 0.044 K, and a line area of 3.52 ± 0.08 K km s^{−1}. The HCO^{+}(3–2) spectrum shows a slight asymmetry with a red wing. The J(4–3) line of HCO^{+} at 356.7 GHz was observed on Mar. 9 with a main-beam temperature of 1.18 ± 0.057 K and a line area of 3.52 ± 0.15 K km s^{−1}. The HCO^{+}(4–3) spectrum shows a redward wing that is offset by 3.6 km/s relative to the main line. The intensity of the red wing is 0.17 ± 0.057 K. The linewidths (FWHM) of the two components are 2.84 ± 0.14 km/s (main line) and 3.17 ± 0.89 km/s (red wing). The CO J(2–1) line at 230.5 GHz was detected on Mar. 5 with a main-beam temperature of 0.27 ± 0.021 K and a line area of 0.67 ± 0.053 K km s^{−1}. The column densities of CO and HCO^{+} are estimated to be 6.1 × 10^{14} cm^{−2} and 2.7 × 10^{12} cm^{−2} (assuming a line-excitation temperature of 40 K and optically thin conditions for the region of emission of CO and HCO^{+}). We calculate the following production rates: CO, 1.7 × 10^{29} molecules/s; HCO^{+}, 7.0 × 10^{26} molecules/s.”

J. E. Wink, Institut de Radio Astronomie Millimètrique (IRAM), Grenoble; D. Bockelée-Morvan, N. Biver, P. Colom, J. Crovisier, E. Gérard, and H. Rauer, Observatoire de Paris; D. Despois, Observatoire de Bordeaux; R. Moreno and G. Paubert, IRAM, Granada; and J. K. Davies and W. R. F. Dent, Joint Astronomy Centre, Hilo, report the detection of SO_{2} in comet C/1995 O1 on Mar. 18.3 UT at the IRAM Plateau de Bure interferometer, using its five antennae in a single-dish mode: “The SO_{2} 11(1,11)–10(0,10) rotational line at 221.965 GHz was detected with an integrated line area on a T_{A}^{∗} scale of 0.08 ± 0.01 K km s^{−1}. Assuming thermal excitation at 80 K, the derived SO_{2} production rate is ∼ 6 × 10^{27} molecules/s. The production rate might be larger if the excitation of the rotational lines of SO_{2} is dominated by radiative excitation of the electronic bands in the ultraviolet. This is the first detection of SO_{2} in a comet, one of the sources of SO (Lis et al., IAUC 6573). Comparison of observations of SO 6(5)–5(4) 219.949-GHz and H_{2}S 2(2,0)–2(1,1) 216.710-GHz lines, made simultaneously in both single-dish and interferometric modes at IRAM on Mar. 13.3, show that the bulk of SO does not come directly from the nucleus, in contrast to H_{2}S.”

1997 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green